OpenStudy (anonymous):

A physical education class has three times as many girls as boys. During a class basketball game, the girls average 18 points each, and the class as a whole averages 17 points per person. How many points does each boy score on average?

OpenStudy (experimentx):

3 times boys = girls or opposite?

OpenStudy (anonymous):

Do you know how to solve it?

OpenStudy (experimentx):

of course

OpenStudy (anonymous):

Can you give me an example about this?

OpenStudy (experimentx):

let a class have times boys than girls if x be girls than the no of boys = 3x now if girls hit on average 18, then total points scored by girls = 18x the total number of students = x +3x = 4x the total points scored = 17*(4x) now total points scored by boys = 17*(4x) - 18x average scored by boys = (17*(4x) - 18x)/3x

OpenStudy (anonymous):

How did you get 4x? Where 4x come from?

OpenStudy (experimentx):

adding girls + boys

OpenStudy (anonymous):

You meant 3x (girls) + 1x (boys), right?

OpenStudy (experimentx):

yes

OpenStudy (anonymous):

68x - 18x/3x?

OpenStudy (experimentx):

68x is the total no of points scored by students, right so points scored by girls = 18x ... isn't it?

OpenStudy (anonymous):

Yes, it is.

OpenStudy (experimentx):

so subtract it form total points to get the points scored by boys, .. and you have points scored by boys and also you know total no of boys, so calculate average

OpenStudy (anonymous):

68x-18x = 50x/3x?

OpenStudy (experimentx):

cut out x's on right had side .. you have 50/3 = 16.666